Lot conforming non conforming
A 88 12

b

c

165

260

35

40

A population of 600 semiconductor wafers contain wafers from three lots. The wafers are categorized by lot and by whether they conform to a thickness specificification with the results shown in the following table.

a.) What is the probability that the wafer is conforming?

b.) What is the probability that the wafer is from Lot A and is conforming?

c.) Given that the wafer is from Lot A, what is the probability that it is conforming?
Answer #1

Ans:

a)

P(conforming)=(88+165+260)/600=0.855

b)

P(lot A and conforming)=88/600=0.147

c)

P(conforming/lot A)=88/(88+12)=0.88